OpenStudy (anonymous):

What are the directrices of the hyperbola given by the equation y^2/49-x^2/9=1

OpenStudy (anonymous):

i am confused because so far as i know a hyperbola does not have a directrix

OpenStudy (anonymous):

it does have two asymptotes however the asymptotes for \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\] are\[y=\pm\frac{a}{b}x\]

OpenStudy (anonymous):

Thank you, that equation actually helps!

OpenStudy (anonymous):

yw